OpenStudy (anonymous):

The following rational function describes blood concentration of a certain drug taken via IV over time, find a. the horizontal or oblique asymptote(s), if any, b. the vertical asymptote(s), if any, c. describe their possible meanings. f(x)= x+1/x

OpenStudy (anonymous):

Ok, one step at the time. First the a) assignment. Definition of horizontal asymptote: A horizontal asymptote is a y-value which a function approaches but does not actually reach. To find it we have to remove everything except biggest exponents in numerator and denominator in given function: \[f(x) = \frac{ x+1 }{ x } \] Biggest exponent of x is x^1 found in both numerator and denominator, so we have: \[f(x) = \frac{ x }{ x }\] Both numerator and denominator have same exponent of x, so they cancel each other out, and we are left with 1/1, which is 1. Horizontal asymptote is 1.

OpenStudy (anonymous):

Thank you so much...... your the best!!!!

OpenStudy (anonymous):

I might need more help.. I hope you don't mind......

OpenStudy (anonymous):

sure, one step at the time;) feel free to ask if you don't understand any part of this solution. here is a good reference for horizontal asymptotes if you need more info: http://www.freemathhelp.com/finding-horizontal-asymptotes.html

OpenStudy (anonymous):

now the b) part Definition of vertical asymptote: Vertical asymptotes are vertical lines which correspond to the zeroes of the denominator of a rational function. As you can see from the definition, we have to find values when will the denominator of our given function be zero in order to find vertical asymptotes. \[f(x) = \frac{ x+1 }{ x }\] Denominator is x, and it will be zero when x=0. Hence, vertical asymptote is 0.